please help. A boy on the bridge throws a stone horizontally with a speed of 25 m/s releasing the stone from a point 19.6m above the surface of the river. How far from the point directly below the boy will the stone strike the water?
This is projectile right? How to solve this? Thank you.

Answer by Alex.33(110)

Me:
Ah this should be a problem posted on physics.com...
Checked physics.com
Nope
Better answer it here, though it should be categorized as "Classic Physics"
=======================================================
Regardless of what's happenning horizontally, vertically its initial speed is 0. It has the accelaration(gravity) of roughly 9.8 m/s^2
The vertical displacement: y=gt^2/2
y=19.6 m(since totally it goes vertically 19.6 metres), g=9.8 m/s^2. Solve. t=2 s.
So the time from start to hitting water is 2 s. This won't change regardless of what's going on horizontally due to the natural properties of these vectors.
Go back. Horizontally it travelled x=vt=2*25=50 metres.
Pretty far!
=======================================================
Anyway hope that helps! And not to be misleading, feel free to ask any physics qustions here, because they are, fundamentally, applications of math, and algebra.